# Repetita Iuvant: Data Repetition Allows SGD to Learn High-Dimensional Multi-Index Functions

### v1

@@ 1 Introduction
Gradient Descent-based algorithms such as Stochastic Gradient Descent (SGD) and its variations, are a fundamental tool in training neural networks, and are therefore the subject of intense theoretical scrutiny. Recent years have witnessed significant advancements in understanding their dynamics and the learning mechanisms of neural nets. Significant progress has been made, in particular, in the case of two-layer networks thanks, in part, to the so-called mean-field analysis (Mei et al., 2018; Chizat and Bach, 2018; Rotskoff and Vanden-Eijnden, 2022; Sirignano and Spiliopoulos, 2020)). A large part of the theoretical approach focused on one-pass optimization algorithms, where each iteration involves a new fresh batch of data. In particular, in the mathematically solvable case of high-dimensional synthetic Gaussian data, and a low dimensional a multi-index target function model, the class of functions efficiently learned by these one-pass methods has been thoroughly analyzed in a series of recent works, and have been shown to be limited by the so-called information exponent (Ben Arous et al., 2021) and leap exponent (Abbe et al., 2022, 2023). These analyses have sparked many follow-up theoretical works over the last few months, see, e.g. Damian et al. (2022, 2023); Dandi et al. (2023); Bietti et al. (2023); Ba et al. (2023); Moniri et al. (2023); Mousavi-Hosseini et al. (2023); Zweig and Bruna (2023), leading to a picture where SGD dynamics is governed by the information exponent, that is, the order of the first non-zero Hermite coefficient of the target.
@@
There are good reason to believe this to be the case. Indeed, it has been observed very recently (Dandi et al., 2024) that iterating twice over large ($O(d)$) datasets was enough to alter the picture reached in Ben Arous et al. (2021); Abbe et al. (2022, 2023). Indeed Dandi et al. (2024) showed that their exist functions that can be learned over few (just two) batch repetition, while they require a much larger (i.e. polynomial in $d$) number of steps otherwise. While the set of function considered in this work was limited, the conclusion is indeed surprising, thus motivating the following question:
@@
We consider an analytically tractable a single-pass training scheme to model these changes: we process one Gaussian sample at a time, but the gradient step is repeated twice! This simple adjustment to the idealized SGD optimization scheme, which is actually often used in machine learning in Lookahead optimizers (Zhang et al., 2019), Extragradient methods (Korpelevič, 1976), or in the context of Sharpness Aware Minimization (SAM Foret et al. (2021)), will be seen to make single-pass algorithms drastically more efficient, and, as we will see, as efficient than our best algorithms in most cases.
We show in particular that most multi-index function are learned with total sample complexity either $O(d)$ or $O(d\log d)$ if they are related to the presence of a symmetry, without any need for preprocessing of the data that is performed in various works (See e.g., Mondelli and Montanari (2018); Luo et al. (2019); Maillard et al. (2020); Chen and Meka (2020)). Our results thus demonstrate that shallow neural networks turns out to be significantly more powerful than previously believed (Ben Arous et al., 2021; Abbe et al., 2022, 2023; Arnaboldi et al., 2023) in identifying relevant feature in the data. In fact, they are shown in many (but not all) cases to saturate the bounds predicted by statistical queries (SQ) (Damian et al., 2024) (rather than the more limited correlation statistical queries (CSQ) bounds that were characteristic of the former consensus (Ben Arous et al., 2021; Abbe et al., 2023)). Our conclusions follow from the rigorous analysis of the overlaps between the trained weight vectors with the low-dimensional relevant subspace that draws inspiration from the pivotal works of Saad and Solla (1995); Coolen and Saad (2000); Coolen et al. (2000); Ben Arous et al. (2021).
@@ 2 Setting and Main Contributions
Given a dataset of $N$ labeled examples $\{\bm{z}^{\nu},y^{\nu}\}_{\nu\in[N]}$ in $d$ dimensions, we analyze the learning properties of fully connected two-layer networks with first layer weights $W=\{\bm{w}_{j}\}_{j\in[p]}$, second layer $\bm{a}\in\mathbb{R}^{p}$, and activation function $\sigma:\mathbb{R}\to\mathbb{R}$
$f(\bm{z};W,\bm{a})=\frac{1}{p}\sum_{j=1}^{p}a_{j}\sigma\left(\langle\bm{z},\bm{w}_{j}\rangle\right)$ (1)
We consider target functions that are dependent only on few orthogonal relevant directions $k=O_{d}(1)$ encoded in the target’s weight matrix $W^{\star}=\{\bm{w}^{\star}_{r}\}_{r\in[k]}\in\mathbb{R}^{k\times d}$:
$y^{\nu}=f^{\star}(\bm{z}^{\nu})=h^{\star}(W^{\star}\bm{z}^{\nu})$ (2)
This model for structured data is usually referred to as a multi-index model. Our main objective will be to analyze how efficiently two-layer nets adapt to this low-dimensional structure during training.
@@
$\mathcal{R}(W,\bm{a})=\mathbb{E}_{\bm{z},y}\left[\mathcal{L}(f(\bm{z},W,\bm{a}),y)\right]$ (3)
@@
This objective is non-convex, and we do not have direct access to the expectation over the data. A central question in Machine Learning is therefore to quantify the properties of the weights attained at convergence by different algorithms. In this manuscript we consider a specific class of one-pass training routines to minimize the empirical risk. First, we partition the dataset in disjoint mini-batches of size $n_{b}$, i.e. $\mathcal{D}=\cup_{t\in[T]}\{\bm{x}^{\nu},y^{\nu}\}_{\nu\in[tn_{b},(t+1)n_{b}]}$ where $T$ is the total number of iterations considered. For each algorithmic step, we update the hidden layer neurons $W=\{\bm{w}_{j}\}_{j\in[p]}$ with the program $\mathcal{A}:(\bm{w}_{j,t},\gamma,\rho,n_{b})\in\mathbb{R}^{d+3}\to\bm{w}_{t+1,j}\in\mathbb{R}^{d}$.
@@
$\bm{w}_{t+1,j}$ $=\mathcal{A}(\bm{w}_{t},\gamma,\rho,n_{b})$ (4)
$=\bm{w}_{t,j}-\frac{\gamma}{n_{b}}\sum_{\nu\in[n_{b}]}\nabla_{\bm{w}_{t,j}}\mathcal{L}\left(f(\bm{z}^{\nu};\tilde{W}_{t}(\rho),\bm{a}_{0}),y^{\nu}\right)$
$\tilde{\bm{w}}_{t,j}(\rho)$ $=\bm{w}_{t,j}-\rho\nabla_{\bm{w}_{t,j}}\mathcal{L}\left(f(\bm{z}^{\nu};W_{t},\bm{a}_{0}),y^{\nu}\right)$
This class of algorithms defined by the optimizer step $\mathcal{A}(\cdot\,,\gamma,\rho,n_{b})$, in which we consider the final gradient update after taking a linear combination of the current iterate and its current gradient (noted as $\tilde{W}_{t}(\rho)$ in the above), is broadly used in different contexts. Indeed, routines with positive $\rho$ parameters correspond to Extragradient methods (EgD) (Korpelevič, 1976), while with negative $\rho$ have been recently used in the context of Sharpness Aware Minimization (SAM) (Foret et al., 2021). For clarity, we present our theoretical results with the above, easily interpretable, Algorithm 1. However, our theoretical claims are valid in a more general setting; we refer to Appendices (A, B) for additional results on more general optimizer steps.
As previously stated, the central object of our analysis is the efficiency of the network to adapt to the low-dimensional structure identified by $W^{\star}$. Therefore, we focus on the learned representations by the first layer weights $W_{t}$ while keeping the second layer weights $\bm{a}_{t}=\bm{a}_{0}$ fixed during training. This assumption is favorable to performing the theoretical analysis and is largely used in the theoretical community (e.g., Damian et al. (2022); Ba et al. (2022); Bietti et al. (2022)).
Numerous theoretical efforts have been devoted to understanding the sample complexity needed to learn multi-index targets. However, an important aspect of many algorithmic routines considered is to exploit a clever warm start to initialize the iterates (Mondelli and Montanari, 2018; Luo et al., 2019; Maillard et al., 2020; Chen and Meka, 2020). We show in this work that such preprocessing of the data is not needed to learn all polynomial multi-index functions in $T=O(d\log d)$ with the routine in Alg. 1. As a way to show this, we consider the extreme case in which the initial iterates $W_{0}$ lives in the orthogonal subspace of the target’s weights $W^{\star}$, i.e., the ‘coldest’ possible start.
@@
The essential object in our analysis is the evolution of the correlation between the hidden neurons $W=\{\bm{w}_{j}\}_{j\in[p]}$ and the target’s weights $W^{\star}$, as a function of the algorithmic time steps. More precisely, we are interested in the number of algorithmic steps needed to achieve an order one correlation with the target’s weights $W^{\star}$, known as weak recovery.
@@
$V^{\star}={\rm span}(\bm{w}^{\star}_{1},\dots,\bm{w}^{\star}_{k})$ (5)
We define the following weak recovery stopping time for a parameter $\eta\in(0,1)$ independent from $d$:
@@
Recent works have sharply theoretically characterized the weak recovery stopping time for a variety of multi-index targets learned with one-pass Stochastic Gradient Descent. First Ben Arous et al. (2021) for the single-index case, then Abbe et al. (2022) for the multi-index one have unveiled the presence of an Information Exponent $\ell$ characterizing the time complexity needed to weakly recover the target subspace $V^{\star}$, defined in the single-index case as
@@
These bounds have been improved by Damian et al. (2023) up to order $d^{\ell/2}$; the latter matches the so-called Correlated Statistical Query lower bound, which considers queries of the form $\mathbb{E}\left[y\phi(\bm{z})\right]$.
However, this Information Exponent is not the right measure of complexity for low-dimensional weak recovery. Indeed, Damian et al. (2024) introduce a new Generative Exponent $\ell^{\star}$ (generative exponent) governing the weak recovery time scaling for algorithms in the Statistical Query (SQ) or Low Degree Polynomial (LDP) family. This exponent is defined as
@@
which allows for the application of arbitrary transformations of the labels before performing the queries. Under this definition, the “hard” problems with $\ell^{\star}>2$ are also impossible to learn for the best first-order algorithm, Approximate Message Passing (Barbier et al., 2019; Celentano et al., 2021; Mondelli and Montanari, 2018; Luo et al., 2019; Maillard et al., 2020).
@@
When the optimal transformation $f$ is known, a simple SGD algorithm on $(\bm{z},f(y))$ achieves the Generative Exponent lower bound. However, one problem remains: does there exist a class of SGD-like algorithms that can achieve this lower bound agnostically?
@@
Although the performance of single-pass SGD is limited by the Information Exponent of $h^{\star}$, the situation drastically changes when multiple-pass algorithms are considered. Recently, Dandi et al. (2024) proved that non-even single-index targets are weakly recovered in $T=O(d)$ when considering extensive batch sizes with multiple pass. This begs the question of how good can a “reusing” algorithm be. We answer this question for the class of polynomial link functions:
@@
This shows that a simple gradient-based algorithm, which only requires seeing the data twice, is optimal for learning any polynomials.
@@
The ability of learning relevant features/representations in the data is arguably one of the key properties of neural networks. However, a considerable attention in the theoretical community has been devoted to understanding the lazy training of two-layer networks (Jacot et al., 2018) where features are not effectively learned. Indeed, in this regime, shallow networks are equivalent to kernel machines, a class of fixed feature methods which are not able to adapt efficiently to low-dimensional relevant structure in the data (see e.g. Ghorbani et al. (2020)). Thus, understanding how representations are learned and providing corrections to the kernel regime is a central question in machine learning theory and has been explored in a variety of works (see e.g., Dudeja and Hsu (2018); Atanasov et al. (2022); Bietti et al. (2022); Damian et al. (2022); Petrini et al. (2022)). In this manuscript we provably characterize the number of iterations needed from Algorithm 1 to learn the relevant features of the data, i.e., attain weak recovery of the target subspace. Our results provide the scaling of the relevant hyperparameters $(\gamma,\rho)$ with the diverging input dimension $d$. The minibatch size will be fixed to $n_{b}=1$ in the main body, and we refer to Appendix B for the extension to larger batch sizes.
@@
  • We prove for single-index targets that one-pass gradient-based algorithms are able to surpass the Correlation Statistical Query (CSQ) limitations established by the Information Exponent (Ben Arous et al., 2021).
  • We unveil that the dynamics of the family of Algorithm 1 are governed by the generative exponent introduced in Damian et al. (2024) with an additional polynomial restriction on the transformation of the output.
  • We generalize the notion of generative exponent to multi-index targets. Similarly to the single-index case, one-pass gradient-based algorithms can overcome CSQ performance dictated by the Leap Coefficient (Abbe et al., 2022).
@@
  • We validate the formal theoretical claims with detailed numerical illustrations. The code to reproduce representative figures is available in the Github repository https://github.com/IdePHICS/Repetita-Iuvant.

### v2

@@ 1 Introduction
Gradient Descent-based algorithms such as Stochastic Gradient Descent (SGD) and its variations, are a fundamental tool in training neural networks, and are therefore the subject of intense theoretical scrutiny. Recent years have witnessed significant advancements in understanding their dynamics and the learning mechanisms of neural nets. Significant progress has been made, in particular, in the case of two-layer networks thanks, in part, to the so-called mean-field analysis (Mei et al., 2018; Chizat and Bach, 2018; Rotskoff and Vanden-Eijnden, 2022; Sirignano and Spiliopoulos, 2020)). A large part of the theoretical approach focused on one-pass optimization algorithms, where each iteration involves a new fresh batch of data. In particular, in the mathematically solvable case of high-dimensional synthetic Gaussian data, and a low dimensional a multi-index target function model, the class of functions efficiently learned by these one-pass methods has been thoroughly analyzed in a series of recent works, and have been shown to be limited by the so-called information exponent (Ben Arous et al., 2021) and leap exponent (Abbe et al., 2022, 2023). These analyses have sparked many follow-up theoretical works over the last few months, see, e.g. Damian et al. (2022, 2023); Dandi et al. (2024a); Bietti et al. (2023); Ba et al. (2023); Moniri et al. (2024); Mousavi-Hosseini et al. (2023); Zweig and Bruna (2024), leading to a picture where SGD dynamics is governed by the information exponent, that is, the order of the first non-zero Hermite coefficient of the target.
@@
There are good reason to believe this to be the case. Indeed, it has been observed very recently (Dandi et al., 2024b) that iterating twice over large ($O(d)$) datasets was enough to alter the picture reached in Ben Arous et al. (2021); Abbe et al. (2022, 2023). Indeed Dandi et al. (2024b) showed that their exist functions that can be learned over few (just two) batch repetition, while they require a much larger (i.e. polynomial in $d$) number of steps otherwise. While the set of function considered in this work was limited, the conclusion is indeed surprising, thus motivating the following question:
@@
We consider a simple yet analytically tractable single-pass training scheme: processing one Gaussian sample at a time, but repeating the gradient step twice. This minor adjustment, common in Lookahead optimizers (Zhang et al., 2019), Extragradient methods (Korpelevič, 1976), and Sharpness-Aware Minimization (SAM) (Foret et al., 2021), significantly boosts efficiency. In fact, we will see that this approach matches the performance of our best algorithms in most cases.
We show in particular that most multi-index function are learned with total sample complexity either $O(d)$ or $O(d\log d)$ if they are related to the presence of a symmetry, without any need for preprocessing of the data that is performed in various works (See e.g., Mondelli and Montanari (2018); Luo et al. (2019); Maillard et al. (2020); Chen and Meka (2020)). Our results thus demonstrate that shallow neural networks turns out to be significantly more powerful than previously believed (Ben Arous et al., 2021; Abbe et al., 2022, 2023; Arnaboldi et al., 2023) in identifying relevant feature in the data. In fact, they are shown in many (but not all) cases to saturate the bounds predicted by statistical queries (SQ) (Damian et al., 2024) (rather than the more limited correlation statistical queries (CSQ) bounds that were characteristic of the former consensus (Ben Arous et al., 2021; Abbe et al., 2023)). Equivalently, it roughly mirrors the prediction of the approximate message passing analysis in (Troiani et al., 2024). Our conclusions follow from the rigorous analysis of the overlaps between the trained weight vectors with the low-dimensional relevant subspace that draws inspiration from the pivotal works of Saad and Solla (1995); Coolen and Saad (2000); Coolen et al. (2000); Ben Arous et al. (2021).
Concurrently with the present work, (Lee et al., 2024) also studies the properties of SGD under data repetition in neural networks. We discuss the differences between our works in Section 3.2.
@@ 2 Setting and Main Contributions
Given a dataset of $N$ labeled examples $\{\mathbf{z}^{\nu},y^{\nu}\}_{\nu\in[N]}$ in $d$ dimensions, we analyze the learning properties of fully connected two-layer networks with first layer weights $W=\{\mathbf{w}_{j}\}_{j\in[p]}$, second layer $\mathbf{a}\in\mathbb{R}^{p}$, and activation function $\sigma:\mathbb{R}\to\mathbb{R}$
$f(\mathbf{z};W,\mathbf{a})=\frac{1}{p}\sum_{j=1}^{p}a_{j}\sigma\left(\langle\mathbf{z},\mathbf{w}_{j}\rangle\right)$ (1)
We consider target functions that are dependent only on few orthogonal relevant directions $k=O_{d}(1)$ encoded in the target’s weight matrix $W^{\star}=\{\mathbf{w}^{\star}_{r}\}_{r\in[k]}\in\mathbb{R}^{k\times d}$:
$y^{\nu}=f^{\star}(\mathbf{z}^{\nu})=h^{\star}(W^{\star}\mathbf{z}^{\nu})$ (2)
This model for structured data is usually referred to as a multi-index model. Our main objective will be to analyze how efficiently two-layer nets adapt to this low-dimensional structure during training. To ensure that all necessary Gaussian expectations exist, we make the following non-restrictive assumption on $\sigma,h^{\star}$:
Assumption 1.
The functions $\sigma$ and $h^{\star}$ have at most polynomial growth, in the sense that there exists a constant $C>0$ such that for all $x\in\mathbb{R},\mathbf{x}\in\mathbb{R}^{k}$,
$|\sigma(x)|\leq C(1+|x|)^{C}\quad\text{and}\quad|h^{\star}(\mathbf{x})|\leq C(1+\|\mathbf{x}\|)^{C}.$ (3)
@@
$\mathcal{R}(W,\mathbf{a})=\mathbb{E}_{\mathbf{z},y}\left[\mathcal{L}(f(\mathbf{z},W,\mathbf{a}),y)\right]$ (4)
@@
This objective is non-convex, and we do not have direct access to the expectation over the data. A central question in Machine Learning is therefore to quantify the properties of the weights attained at convergence by different algorithms. In this manuscript we consider a specific class of one-pass training routines to minimize the empirical risk. First, we partition the dataset in disjoint mini-batches of size $n_{b}$, i.e. $\mathcal{D}=\cup_{t\in[T]}\{\mathbf{x}^{\nu},y^{\nu}\}_{\nu\in[tn_{b},(t+1)n_{b}]}$ where $T$ is the total number of iterations considered. For each algorithmic step, we update the hidden layer neurons $W=\{\mathbf{w}_{j}\}_{j\in[p]}$ with the program $\mathcal{A}:(\mathbf{w}_{j,t},\gamma,\rho,n_{b})\in\mathbb{R}^{d+3}\to\mathbf{w}_{t+1,j}\in\mathbb{R}^{d}$.
@@
$\mathbf{w}_{t+1,j}$ $=\mathcal{A}(\mathbf{w}_{t},\gamma,\rho,n_{b})=\mathbf{w}_{t,j}-\frac{\gamma}{n_{b}}\sum_{\nu\in[n_{b}]}\nabla_{\mathbf{w}_{t,j}}\mathcal{L}\left(f(\mathbf{z}^{\nu};\tilde{W}_{t}(\rho),\mathbf{a}_{0}),y^{\nu}\right)$ (5)
$\tilde{\mathbf{w}}_{t,j}(\rho)$ $=\mathbf{w}_{t,j}-\rho\nabla_{\mathbf{w}_{t,j}}\mathcal{L}\left(f(\mathbf{z}^{\nu};W_{t},\mathbf{a}_{0}),y^{\nu}\right)$
This class of algorithms defined by the optimizer step $\mathcal{A}(\cdot\,,\gamma,\rho,n_{b})$, in which we consider the final gradient update after taking a linear combination of the current iterate and its current gradient (noted as $\tilde{W}_{t}(\rho)$ in the above), is broadly used in different contexts. Indeed, routines with positive $\rho$ parameters correspond to Extragradient methods (EgD) (Korpelevič, 1976), while with negative $\rho$ have been recently used in the context of Sharpness Aware Minimization (SAM) (Foret et al., 2021). For clarity, we present our theoretical results with the above, easily interpretable, Algorithm 1. However, our theoretical claims are valid in a more general setting; we refer to Appendices (B, D) for additional results on more general optimizer steps.
As previously stated, the central object of our analysis is the efficiency of the network to adapt to the low-dimensional structure identified by $W^{\star}$. Therefore, we focus on the learned representations by the first layer weights $W_{t}$ while keeping the second layer weights $\mathbf{a}_{t}=\mathbf{a}_{0}$ fixed during training. This assumption is favorable to performing the theoretical analysis and is largely used in the theoretical community (e.g., Damian et al. (2022); Ba et al. (2022); Bietti et al. (2022)).
Numerous theoretical efforts have been devoted to understanding the sample complexity needed to learn multi-index targets. However, an important aspect of many algorithmic routines considered is to exploit a clever warm start to initialize the iterates (Mondelli and Montanari, 2018; Luo et al., 2019; Maillard et al., 2020; Chen and Meka, 2020). We show in this work that such preprocessing of the data is not needed to learn all polynomial multi-index functions in $T=O(d\log d)$ with the routine in Alg. 1. Instead, we consider a fully agnostic initialization of the network:
Assumption 2.
For each $j\in[p]$, the initial first-layer weight $\mathbf{w}_{0,j}$ is initialized according to the uniform measure on the sphere:
$\mathbf{w}_{0,j}\overset{i.i.d}{\sim}\operatorname{Unif}(\mathbb{S}^{d-1}).$
The second-layer weights are drawn i.i.d in $\{-1,1\}$:
$a_{0,j}\overset{i.i.d}{\sim}\operatorname{Unif}(\{-1,1\}).$
@@
The essential object in our analysis is the evolution of the correlation between the hidden neurons $W=\{\mathbf{w}_{j}\}_{j\in[p]}$ and the target’s weights $W^{\star}$, as a function of the algorithmic time steps. More precisely, we are interested in the number of algorithmic steps needed to achieve an order one correlation with the target’s weights $W^{\star}$, known as weak recovery.
@@
$V^{\star}=\operatorname{span}(\mathbf{w}^{\star}_{1},\dots,\mathbf{w}^{\star}_{k})$ (6)
We define the weak recovery stopping time for a parameter $\eta\in(0,1)$ independent from $d$ as:
@@
Recent works have sharply theoretically characterized the weak recovery stopping time for a variety of multi-index targets learned with one-pass SGD. Ben Arous et al. (2021) for the single-index case, then Abbe et al. (2022) for the multi-index one have unveiled the presence of an Information Exponent $\ell$ characterizing the time complexity needed to weakly recover the target subspace $V^{\star}$, defined in the single-index case as
@@
These bounds have been improved by Damian et al. (2023) up to order $d^{\ell/2}$; the latter matches the so-called Correlated Statistical Query lower bound, which considers queries of the form $\mathbb{E}\left[y\phi(\mathbf{z})\right]$.
However, this Information Exponent is not the right measure of complexity for low-dimensional weak recovery. Some iterative algorithms, for instance, are known to perform drastically better(Barbier et al., 2019; Celentano et al., 2021; Mondelli and Montanari, 2018; Luo et al., 2019; Maillard et al., 2020; Troiani et al., 2024). Recently Damian et al. (2024) introduced a new Generative Exponent $\ell^{\star}$ (generative exponent) governing the weak recovery time scaling for algorithms in the Statistical Query (SQ) or Low Degree Polynomial (LDP) family. This exponent is defined as
@@
which allows for the application of arbitrary transformations of the labels before performing the queries. Under this definition, the “hard” problems with $\ell^{\star}>2$ are also impossible to learn for the best first-order algorithm, Approximate Message Passing (Troiani et al., 2024).
@@
When the optimal transformation $f$ is known, a simple SGD algorithm on $(\mathbf{z},f(y))$ achieves the Generative Exponent lower bound. However, one problem remains: does there exist a class of SGD-like algorithms that can achieve this lower bound agnostically?
@@
Although the performance of single-pass SGD is limited by the Information Exponent of $h^{\star}$, the situation drastically changes when multiple-pass algorithms are considered. Recently, Dandi et al. (2024b) proved that non-even single-index targets are weakly recovered in $T=O(d)$ when considering extensive batch sizes with multiple pass. This begs the question of how good can a “reusing” algorithm be. We answer this question for the class of polynomial link functions:
@@
This shows in particular that a simple gradient-based algorithm, which only requires seeing the data twice, is optimal for learning any polynomials.
@@
Learning relevant features from data is a fundamental property of neural networks. However, much theoretical work has focused on lazy training in two-layer networks (Jacot et al., 2018), where features are not effectively learned. In this regime, shallow networks behave as kernel machines, which struggle to adapt to low-dimensional structures in data (Ghorbani et al., 2020). Understanding representation learning and corrections to the kernel regime is thus a key challenge in machine learning theory (Dudeja and Hsu, 2018; Atanasov et al., 2022; Bietti et al., 2022; Damian et al., 2022; Petrini et al., 2022). In this manuscript we provably characterize the number of iterations needed from Algorithm 1 to learn the relevant features of the data, i.e., attain weak recovery of the target subspace. Our results provide the scaling of the relevant hyperparameters $(\gamma,\rho)$ with the diverging input dimension $d$. The minibatch size will be fixed to $n_{b}=1$ in the main body, and we refer to Appendix D for the extension to larger batch sizes.
@@
  • We prove for single-index targets that one-pass gradient-based algorithms surpass the Correlation Statistical Query (CSQ) limits established by the Information Exponent (Ben Arous et al., 2021).
  • We unveil that the dynamics of the family of Alg. 1 are governed by the generative exponent Damian et al. (2024), with an additional polynomial restriction on the transformation of the output.
  • We generalize the notion of generative exponent to multi-index targets. As in the single-index case, one-pass gradient-based algorithms can overcome CSQ performance dictated by the Leap Coefficient (Abbe et al., 2022). Our definition is consistent with the one of (Troiani et al., 2024).
@@
  • We validate the formal theoretical claims with detailed numerical illustrations. The code to reproduce representative figures is available at https://github.com/IdePHICS/Repetita-Iuvant.